# Language Model Preference Evaluation with Multiple Weak Evaluators

### v1

@@ 1 Introduction
Large Language Models (LLMs) have rapidly transformed various fields within artificial intelligence, particularly natural language processing (NLP) and decision-making systems (Wu et al., 2023; Li et al., 2023a). Despite the remarkable success of LLMs, the need for effective evaluation methods becomes paramount (Liu et al., 2023; Desmond et al., 2024; Siska et al., 2024). Preference evaluation, as one of the most important assessment methods, plays an indispensable role in evaluating and optimizing model performance (Rafailov et al., 2024; Yuan et al., 2024; Dubois et al., 2024b). Existing works usually leverage a powerful LLM (e.g., GPT4 (Achiam et al., 2023)) as the judge for comparing LLMs’ output pairwisely (Li et al., 2023b; Chen et al., 2023; Li et al., 2023b; Wang et al., 2022).
Figure 1: The figure is divided into two parts. The upper part shows the construction of a preference graph from pairwise LLM evaluations, with cycles indicating inconsistent preferences highlighted in red. The lower part presents the evaluation of ten responses for HumanEval (Chen et al., 2021) and MATH (Hendrycks et al., 2021) tasks, generated by Llama3-70B (AI@Meta, 2024). These were evaluated by GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al., 2023), Qwen2-72B (Yang et al., 2024a), and Llama3-8B, showing that even advanced LLMs exhibit significant noise in their evaluations, illustrating their limitations as reliable evaluators.
However, while such model-based pairwise preference evaluations offer a flexible approach, they can lead to contradictory evaluations in the assessment process (Zhang et al., 2024b). For example, an LLM might evaluate three responses and conclude that Response A is better than Response B (A $\succ$ B), Response B is better than Response C (B $\succ$ C), yet paradoxically also rank Response C as better than Response A (C $\succ$ A). These cyclic patterns introduce inconsistencies that undermine the reliability of the evaluation results. We model this conflicting preference via the preference graph. Specifically, a preference graph is constructed with each response as a node and directed edges indicating pairwise preferences—an edge from node A to node B shows that the evaluator preferred response A over B. The noise illustrated by cycles (A $\succ$ B $\succ$ C $\succ$ A) manifests as loops in the preference graph. The whole process we show in the upper part of Figure 1. Ideally, a preference graph should be structured as a directed acyclic graph (DAG) to maintain consistency. As shown in the bottom half of Figure  1, even advanced LLMs like GPT-4-o exhibit significant noise in preference evaluation, highlighting their limitations as weak evaluators.
@@ 2.2 GED: Preference Graph Ensemble and Denoise
Finding a minimum FAS in general is known to be an NP-complete problem, whose computational complexity can be exponential (Karp, 2010; Bodlaender et al., 2012). Therefore, in our experiment, we apply the well-established approximation algorithm proposed in Eades et al. (1993). Details can be found in Appendix D.
@@ 2.3 Applications
In the response selection task, a model $\mathcal{M}$ generates $n$ candidate answers $\{ans_{1},\ldots,ans_{n}\}$ for each question $q\in Q$, with the objective of identifying the optimal answer $ans_{q}^{*}$ for each query. To achieve this, we employ multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$, who provide pairwise preferences among the candidate answers. For each question $q$, we construct a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{q},A_{a},w_{a})$ encapsulates the preferences of evaluator $a$. The vertex set $V_{q}=\{v_{1},\ldots,v_{n}\}$ corresponds to the candidate answers, while the directed arcs $A_{a}$ indicate the pairwise preferences among these responses. Each arc is weighted by $w_{a}$, reflecting the strength of the preference indicated by the evaluator. The construction of preference graphs involves evaluating each pair of candidate answers $ans_{i}$ and $ans_{j}$. Evaluators assess the quality of these answers, assigning a preference that is denoted by a directed arc $(v_{i},v_{j})$ in $A_{a}$, with a corresponding weight $w_{a}(v_{i},v_{j})$ based on the strength of preference. This process is detailed in Appendix B. After collecting the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ for a question $q$, we apply GED to aggregate these graphs. The aggregation begins with merging all preference graphs into a unified graph $G_{q}=(V_{q},A_{q},w_{q})$, which is subsequently processed to remove cycles, resulting in a DAG. From this DAG, we derive the ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$ of the candidate answers, where the highest-ranked answer is selected as $ans_{q}^{*}$. This process is repeated for each question $q\in Q$, yielding a final set of selected answers $ans^{*}=\{ans_{1}^{*},\ldots,ans_{t}^{*}\}$. This set reflects a consensus from multiple evaluators in $\mathcal{A}$, ensuring the chosen answers represent the highest quality responses based on rigorous preference evaluation.
@@
In model ranking task, the goal is to rank a set of models $M=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ based on their responses to a series of questions $Q=\{q_{1},\ldots,q_{t}\}$. A group of evaluators, denoted as $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$, assesses the model outputs by providing preferences for pairs of responses for each question. For each question $q\in Q$, the evaluators generate preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{q},A_{a},w_{a})$ encapsulates the preferences of evaluator $a$ over the models. The vertex set $V_{q}=\{v_{1},\ldots,v_{n}\}$ corresponds to the models, while the directed arcs $A_{a}$ indicate pairwise preferences, with weights $w_{a}:A_{a}\rightarrow\mathbb{R}^{+}$ reflecting the strength of these preferences. The preference graph for a given question $q$ is constructed by evaluating each pair of models $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, represented by nodes $v_{i}$ and $v_{j}$. For their respective answers $ans_{i}=\mathcal{M}_{i}(q)$ and $ans_{j}=\mathcal{M}_{j}(q)$, evaluators provide a preference indicating which answer is favored. This preference is represented by a directed arc $(v_{i},v_{j})$ in $A_{a}$, assigned a weight $w_{a}(v_{i},v_{j})$ based on preference strength. The detailed procedure is outlined in Appendix B. Once the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ are collected for a question $q$, we employ GED to aggregate these graphs. The method begins by merging all preference graphs into a single graph $G_{q}=(V_{q},A_{q},w_{q})$, which is then transformed into a DAG by removing cycles. The final ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$ is derived from this DAG. This process is repeated for each question $q\in Q$, yielding a set of rankings $\{\mathcal{R}_{q}:q\in Q\}$. To compute the overall ranking $\mathcal{R}^{*}$ of the models across all questions, we conduct a ranking ensemble on the set $\{\mathcal{R}_{q}:q\in Q\}$, as detailed in Appendix E. This approach culminates in a final ranking that reflects the models’ performance as assessed by multiple evaluators in $\mathcal{A}$.
@@ 3 Theoretical Analysis
The full proof is given in Appendix C. From the theorem, we can see that the probability of failure decreases exponentially as the number of samples $N$ increases. Meanwhile, this guarantee only requires $\delta_{1}<0.5$ and does not place restrictions on $\delta_{2}$, which are very mild conditions.
@@ 4 Experiments on Response Selection
Llama3-8B 43.90 27.29 22.08 56.67 6.78 31.14
Mistral-7B 23.17 11.80 23.25 39.83 7.03 15.21
Qwen2-7B 48.58 25.71 59.92 76.75 7.70 43.17
Qwen2-72B 57.93 29.58 72.75 84.67 11.52 51.91
@@
Single model Self-consistency 60.98 29.33 73.58 84.91 8.86 51.13
Llama3-8B 62.19 29.31 74.27 83.16 11.31 52.87
with graph denoising 64.02 30.18 74.73 86.00 11.72 53.53
Mistral-7B 67.24 27.70 74.41 83.83 10.50 52.93
with graph denoising 68.73 29.93 74.77 83.91 10.74 53.21
Qwen2-7B 61.58 28.69 74.50 85.41 11.11 52.84
@@
Qwen2-72B 60.97 31.04 74.73 86.47 12.14 53.87
Single evaluator with graph denoising 68.90 31.17 75.33 87.45 12.26 55.42
GED (w/o denoising) 69.25 30.98 74.29 87.17 12.68 54.83
Multiple evaluator GED 70.73 32.44 75.58 88.18 13.33 56.41
@@
In this section, we evaluate the performance of GED on five benchmarks: HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), and GAIA (Mialon et al., 2023). The Qwen2-72B (Yang et al., 2024a) model ($\mathcal{M}$) generates ten candidate responses per question, and we assess the effectiveness of different methods in selecting the best response. For further implementation details, see Appendix A. We evaluate performance using three setups. First, in the single model setting, the baselines include ContraSolver(Zhang et al., 2024b), Self-consistency(Wang et al., 2022), and direct evaluation with models (Llama3-8B, Mistral-7B, Qwen2-7B and Qwen2-72B). Additionally, we include a baseline called ListPreference, where instead of pairwise comparisons, all candidate responses are input into Qwen2-72B for selecting the most appropriate response. Then, in the single evaluator setting, individual evaluators (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) select the best response from $\mathcal{M}$’s outputs, with and without applying GED’s graph denoising. Finally, in the multiple weak evaluators setup, we combine three weaker evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with GED. We present the results of GED and its variant (w/o denoising), which ensembles the preference graphs without the denoising step.
@@
Table 1 presents the results of the response selection task across five benchmarks. GED consistently outperforms baseline methods, including both single model evaluations (single model) and direct response selection by individual models (single evaluator). This demonstrates the strength of aggregating weak evaluators with GED, particularly when coupled with graph denoising, which enhances response quality by filtering out noise and biases. This highlights the effectiveness of aggregating weak evaluators and applying graph denoising to improve response quality. Furthermore, by combining preference graphs from weaker models (Llama3-8B, Mistral-7B, Qwen2-7B), GED surpasses the performance of a much stronger evaluator (Qwen2-72B). This underscores the value of ensemble methods in mitigating the limitations of individual evaluators. Then, the denoising process proves to be crucial for improving consistency and overall response quality. The substantial performance gains observed when using GED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection. Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text. Lastly, to further evaluate GED, we compared its performance with GPT-3.5, GPT-4-o-mini, and GPT-4-o. Due to computational and API cost constraints, we limited the evaluation to 100 data points for each task. As shown in Figure 3, GED consistently outperformed GPT-3.5 across all tasks and surpassed GPT-4-o-mini on challenging benchmarks like HumanEval and GSM8k. These results highlight the superiority of GED, particularly in leveraging multi-weak evaluators and graph denoising to outperform individual state-of-the-art models.
Alabation study.
@@
We evaluate the impact of removing the ensembling step in GED, referred to as the (w/o ensemble) variant. In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix E). For simplicity of presentation, we use Weight Score to represent GED (w/o ensemble) (Weight Score). As shown in Figure 4, all (w/o ensemble) methods consistently underperform compared to GED. This performance gap arises because converting graphs to ranks before aggregation leads to information loss. In contrast, GED ensembles the graphs directly, preserving more detailed preference information and resulting in better final rankings.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have rapidly transformed various fields within artificial intelligence, particularly natural language processing (NLP) and decision-making systems (Wu et al., 2023; Li et al., 2023a). Despite the remarkable success of LLMs, the need for effective evaluation methods becomes paramount (Liu et al., 2023; Desmond et al., 2024; Siska et al., 2024). Preference evaluation, as one of the most important assessment methods, plays an indispensable role in evaluating and optimizing model performance (Rafailov et al., 2024; Yuan et al., 2024; Dubois et al., 2024b). Existing works usually leverage a powerful LLM (e.g., GPT4 (Achiam et al., 2023)) as the judge for comparing LLMs’ output pairwisely (Li et al., 2023b; Chen et al., 2023b; Wang et al., 2022).
Figure 1: The figure is divided into two parts. The upper part shows the construction of a preference graph from pairwise LLM evaluations, with cycles indicating inconsistent preferences highlighted in red. The lower part presents the evaluation of ten responses for HumanEval (Chen et al., 2021) and MATH (Hendrycks et al., 2021) tasks, generated by Llama3-70B (AI@Meta, 2024). These were evaluated by GPT-4-o, GPT-4-o-mini, GPT-3.5 (Achiam et al., 2023), Qwen2-72B (Yang et al., 2024a), and Llama3-8B, revealing a significant cycle rate (details in Appendix A.5) in their evaluations. This highlights the limitations of even advanced LLMs as reliable evaluators.
However, while such model-based pairwise preference evaluations offer a flexible approach, they can lead to contradictory evaluations in the assessment process (Naresh et al., 2024; Zhang et al., 2024b). For example, an LLM might evaluate three responses and conclude that Response A is better than Response B (A $\succ$ B), Response B is better than Response C (B $\succ$ C), yet paradoxically also rank Response C as better than Response A (C $\succ$ A). These cyclic patterns introduce inconsistencies that undermine the reliability of the evaluation results. We model this conflicting preference via the preference graph. Specifically, a preference graph is constructed with each response as a node and directed edges indicating pairwise preferences—an edge from node A to node B shows that the evaluator preferred response A over B. The noise illustrated by cycles (A $\succ$ B $\succ$ C $\succ$ A) manifests as loops in the preference graph. The whole process we show in the upper part of Figure 1. Ideally, a preference graph should be structured as a directed acyclic graph (DAG) to maintain consistency. As shown in the bottom half of Figure  1, even advanced LLMs like GPT-4-o exhibit significant noise in preference evaluation, highlighting their limitations as weak evaluators (Refer to Appendix A.3 for further details).
@@ 2.2 GED: Preference Graph Ensemble and Denoise
Finding a minimum FAS in general is known to be an NP-complete problem, whose computational complexity can be exponential (Karp, 2010; Bodlaender et al., 2012). Therefore, in our experiment, we apply the well-established approximation algorithm proposed in Eades et al. (1993). Details can be found in Appendix L.
@@ 2.3 Applications
In the response selection task, a model $\mathcal{M}$ generates $n$ candidate answers $\{ans_{1},\ldots,ans_{n}\}$ for each question $q\in Q$, with the objective of identifying the optimal answer $ans_{q}^{*}$ for each query. To achieve this, we employ multiple evaluators $\mathcal{A}=\{a_{1},\ldots,a_{k}\}$, who provide pairwise preferences among the candidate answers. For each question $q$, we construct a set of preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{q},A_{a},w_{a})$ encapsulates the preferences of evaluator $a$. The vertex set $V_{q}=\{v_{1},\ldots,v_{n}\}$ corresponds to the candidate answers, while the directed arcs $A_{a}$ indicate the pairwise preferences among these responses. Each arc is weighted by $w_{a}$, reflecting the strength of the preference indicated by the evaluator. The construction of preference graphs involves evaluating each pair of candidate answers $ans_{i}$ and $ans_{j}$. Evaluators assess the quality of these answers, assigning a preference that is denoted by a directed arc $(v_{i},v_{j})$ in $A_{a}$, with a corresponding weight $w_{a}(v_{i},v_{j})$ based on the strength of preference. This process is detailed in Appendix D. After collecting the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ for a question $q$, we apply GED to aggregate these graphs. The aggregation begins with merging all preference graphs into a unified graph $G_{q}=(V_{q},A_{q},w_{q})$, which is subsequently processed to remove cycles, resulting in a DAG. From this DAG, we derive the ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$ of the candidate answers, where the highest-ranked answer is selected as $ans_{q}^{*}$. This process is repeated for each question $q\in Q$, yielding a final set of selected answers $ans^{*}=\{ans_{1}^{*},\ldots,ans_{t}^{*}\}$. This set reflects a consensus from multiple evaluators in $\mathcal{A}$, ensuring the chosen answers represent the highest quality responses based on rigorous preference evaluation.
@@
In model ranking task, the goal is to rank a set of models $M=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ based on their responses to a series of questions $Q=\{q_{1},\ldots,q_{t}\}$. A group of evaluators, denoted as $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$, assesses the model outputs by providing preferences for pairs of responses for each question. For each question $q\in Q$, the evaluators generate preference graphs $\{G_{a}:a\in\mathcal{A}\}$, where each graph $G_{a}=(V_{q},A_{a},w_{a})$ encapsulates the preferences of evaluator $a$ over the models. The vertex set $V_{q}=\{v_{1},\ldots,v_{n}\}$ corresponds to the models, while the directed arcs $A_{a}$ indicate pairwise preferences, with weights $w_{a}:A_{a}\rightarrow\mathbb{R}^{+}$ reflecting the strength of these preferences. The preference graph for a given question $q$ is constructed by evaluating each pair of models $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, represented by nodes $v_{i}$ and $v_{j}$. For their respective answers $ans_{i}=\mathcal{M}_{i}(q)$ and $ans_{j}=\mathcal{M}_{j}(q)$, evaluators provide a preference indicating which answer is favored. This preference is represented by a directed arc $(v_{i},v_{j})$ in $A_{a}$, assigned a weight $w_{a}(v_{i},v_{j})$ based on preference strength. The detailed procedure is outlined in Appendix D. Once the preference graphs $\{G_{a}:a\in\mathcal{A}\}$ are collected for a question $q$, we employ GED to aggregate these graphs. The method begins by merging all preference graphs into a single graph $G_{q}=(V_{q},A_{q},w_{q})$, which is then transformed into a DAG by removing cycles. The final ranking $\mathcal{R}_{q}=\{v_{1}\succ v_{2}\succ\ldots\succ v_{n}\}$ is derived from this DAG. This process is repeated for each question $q\in Q$, yielding a set of rankings $\{\mathcal{R}_{q}:q\in Q\}$. To compute the overall ranking $\mathcal{R}^{*}$ of the models across all questions, we conduct a ranking ensemble on the set $\{\mathcal{R}_{q}:q\in Q\}$, as detailed in Appendix M. This approach culminates in a final ranking that reflects the models’ performance as assessed by multiple evaluators in $\mathcal{A}$.
@@ 3 Theoretical Analysis
The full proof is given in Appendix E. From the theorem, we can see that the probability of failure decreases exponentially as the number of samples $N$ increases. Meanwhile, this guarantee only requires $\delta_{1}<0.5$ and does not place restrictions on $\delta_{2}$, which are very mild conditions.
@@ 4 Experiments on Response Selection
Llama3-8B 43.90 27.29 22.08 56.67 6.78 31.34
Mistral-7B 23.17 11.80 23.25 39.83 7.03 21.01
Qwen2-7B 48.58 25.71 59.92 76.75 7.70 43.73
Qwen2-72B 57.93 29.58 72.75 84.67 11.52 51.29
@@
Single model Self-consistency 60.98 29.33 73.58 84.91 8.86 51.53
Llama3-8B 62.19 29.31 74.27 83.16 11.31 52.04
with graph denoising 64.02 30.18 74.73 86.00 11.72 53.33
Mistral-7B 67.24 27.70 74.41 83.83 10.50 52.73
with graph denoising 68.73 29.93 74.77 83.91 10.74 53.61
Qwen2-7B 61.58 28.69 74.50 85.41 11.11 52.25
@@
Qwen2-72B 60.97 31.04 74.73 86.47 12.14 53.07
Single evaluator with graph denoising 68.90 31.17 75.33 87.45 12.26 55.02
Multi-MV 66.18 29.57 74.77 86.42 11.72 53.73
GED (w/o denoising) 69.25 30.98 74.29 87.17 12.68 54.87
Multiple evaluator GED 70.73 32.44 75.58 88.18 13.33 56.05
@@
In this section, we evaluate the performance of GED on five benchmarks: HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), and GAIA (Mialon et al., 2023). The Qwen2-72B (Yang et al., 2024a) model ($\mathcal{M}$) generates ten candidate responses per question, and we assess the effectiveness of different methods in selecting the best response. For further implementation details, see Appendix A. We evaluate performance using three setups. First, in the single model setting, the baselines include ContraSolver(Zhang et al., 2024b), Self-consistency(Wang et al., 2022), and direct evaluation with models (Llama3-8B, Mistral-7B, Qwen2-7B and Qwen2-72B). Additionally, we include a baseline called ListPreference, where instead of pairwise comparisons, all candidate responses are input into Qwen2-72B for selecting the most appropriate response. Then, in the single evaluator setting, individual evaluators (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) select the best response from $\mathcal{M}$’s outputs, with and without applying GED’s graph denoising. Finally, in the multiple weak evaluators setup, we combine three weaker evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with GED. We also include a new baseline referred to as Multi-MV, which just takes the majority for each pairwise comparison as the preferred answer. We present the results of GED and its variant (w/o denoising), which ensembles the preference graphs without the denoising step.
@@
Table 1 presents the results of the response selection task across five benchmarks. GED consistently outperforms baseline methods, including both single model evaluations (single model) and direct response selection by individual models (single evaluator). This demonstrates the strength of aggregating weak evaluators with GED, particularly when coupled with graph denoising, which enhances response quality by filtering out noise and biases. This highlights the effectiveness of aggregating weak evaluators and applying graph denoising to improve response quality. Furthermore, by combining preference graphs from weaker models (Llama3-8B, Mistral-7B, Qwen2-7B), GED surpasses the performance of a much stronger evaluator (Qwen2-72B). This underscores the value of ensemble methods in mitigating the limitations of individual evaluators. Then, the denoising process proves to be crucial for improving consistency and overall response quality. The substantial performance gains observed when using GED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection. We also evaluated Multi-MV, which aggregates pairwise comparisons using majority voting across evaluators. While Multi-MV offers improvements over individual evaluators, it underperforms compared to GED, highlighting GED’s ability to capture more nuanced evaluation signals and reduce inconsistencies. Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text. Lastly, to further evaluate GED, we compared its performance with GPT-3.5, GPT-4-o-mini, and GPT-4-o. Due to computational and API cost constraints, we limited the evaluation to 100 data points for each task. As shown in Figure 3, GED consistently outperformed GPT-3.5 across all tasks and surpassed GPT-4-o-mini on challenging benchmarks like HumanEval and GSM8k. These results highlight the superiority of GED, particularly in leveraging multi-weak evaluators and graph denoising to outperform individual state-of-the-art models.
Ablation study.
@@
We evaluate the impact of removing the ensembling step in GED, referred to as the (w/o ensemble) variant. In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix M). For simplicity of presentation, we use Weight Score to represent GED (w/o ensemble) (Weight Score). As shown in Figure 4, all (w/o ensemble) methods consistently underperform compared to GED. This performance gap arises because converting graphs to ranks before aggregation leads to information loss. In contrast, GED ensembles the graphs directly, preserving more detailed preference information and resulting in better final rankings.